Xj are Random Variables which are continous and independent, means of Xj are 4 (E[ Xj ]=4), PDFs(Probability density function) of Xj are exponantial and they have the same PDFs (Probability density function). Let N is a RV(random variable) with Poisson PMF(probability mass function), whose mean is 7.(E[N]=7) Y = X1 + … + XN. If you find a function of Y which produces moment from table or by using a software like Matlab, Find PDF(Probability density function), Mean[E] and Variance(Var) of function of Y without process method.
Answer #1

g=lun sedaj XN Given X; ~ Exponential (mean = 4) indeperdita The pay of xj is $(*) where I >o .4 Hiyo otherwise It is a RV wiMean of Y +X en Enx E(Y) = ELĖ (Y/N=n)] Now E(Y/Ar=n) = E(x,+X,+ + Xalatan = E(x, +X ₂ + . nEXI) [Since Xis are independents